OpenStudy (anonymous):

Find constants A and B that make the equation true 2x-9/x^2- x- 6 = A/x-3 + B/x+2 / as in fraction and ^2 as squared???

OpenStudy (lopezking1):

ummmm try posting this in math lol :)

OpenStudy (lopezking1):

@Shelby2124

OpenStudy (anonymous):

Oh sorry I thought this was in the math section... Haha first time using this:p

OpenStudy (anonymous):

@lopezking1

OpenStudy (lopezking1):

Lol it's okay :)

OpenStudy (anonymous):

Any idea? @lopezking1

OpenStudy (lopezking1):

on the question

OpenStudy (anonymous):

Yes @lopezking1

OpenStudy (lopezking1):

like that \[\frac{ 2x-9 }{ x ^{2}-x-6 }=\frac{ a }{ x-3 }+\frac{ b }{ x+2 }\]

OpenStudy (lopezking1):

@Shelby2124

OpenStudy (anonymous):

Thanks @lopezking1